15x + 2 sqrt x - 1 = 0 sqrt = square root of x
Solving equations in quadriatic form
please help if can

Answer by stanbon(75887)

15x + 2 sqrt x - 1 = 0
2sqrtx = 1-15x
Square both sides to get:
4x = 1-30x+225x^2
225x^2-34x+1=0
Solve with quadratic formula:
x=[34+-sqrt(34^2-4*225*1]/[2*225]
x=[34+-sqrt(256)]/450
x=[34+-16]/450
x=50/450 or x= 18/450
x=1/9 or x=1/25
----------------
Check the answers:
15(1/9)+2*(1/3)-1=0
5/3+2/3-1=0
0=0
------------
15(1/25)+2(1/5)-1=0
(3/5)+2/5 - 1=0
0=0
---------------
Cheers,
Stan H.